OpenStudy (anonymous):

need help ,, confused if A = 15 cents an Apple and B = 10 cent Bananas which of the following represents the total value of the purchase in cents ? my options a.)A+B b.)25 (A+B) c.)15A + 10B I would say its b simply because it shows the total value im asking if there's any rule or trick here to make the answer different then the one i chose, thanks

OpenStudy (jdoe0001):

hmmmm, well, did she buy apples only or bananas only? :)

OpenStudy (anonymous):

theres no she, doesn't say how much either just simply the price at which they are purchased

OpenStudy (jdoe0001):

say you are a cashier at a market someone goes to the register to pay for the apples and bananas so say she has 3 apples and 5 bananas how much would you charge her?

OpenStudy (anonymous):

95

OpenStudy (jdoe0001):

95 cents? why? because what you did is ok she has 5 bananas, bananas are 10c each, 10 x B = 50 ok she has 3 apples, apples are 15c each, 15 x A = 45 so you charged her 10B + 15A

OpenStudy (anonymous):

ahhh i see there was no info on how many so 25 doesnt exist thank u

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

algabriacly you have :) are given apples=15 cents A=15 C bannas= 10 cents B=10 C You are asked together what are they in cents? A+B if we plug in then 15C+10C= 25C but something looks fishy i think D works also

OpenStudy (anonymous):

what D timo?

OpenStudy (anonymous):

C sorry lol i remembered the last one as D always

OpenStudy (anonymous):

A+B works also cuzz they can say 3 Bs and 4 As 4A+3B plug in again 4(15C)+3(10C) 60C+30C=90C

OpenStudy (anonymous):

i know a and b would work but i see they gave me two options both 15a and 10 b the other 10 a 15 b u know that half and half rule with multiple choice its either one of those

OpenStudy (anonymous):

B wouldn't work

OpenStudy (anonymous):

it will only work if you have same number of A and B that is for example 1 A and 1 B 2 A and 2 B n A n B It is a sepcial case.

OpenStudy (anonymous):

A=15c*apple B=10c*bannana Therefore total A+B --- 15A+10B is wrong because if you plug in algabriacly you will get 15 and 10 twice

OpenStudy (anonymous):

Another hint i can give is that A=15c*apple B=10c*bannana work because you let apple and bannana vary. THe word problem says 15c an apple, 10c a bannana. You prove it to yourself by saying 15c*apple let apple vary so you know 15c*1 =15c, 15c*2=30c so your expression works out. same applies to bannana 10c*1=10, 20c*2=20c